# In2Core: Leveraging Influence Functions for Coreset Selection in Instruction Finetuning of Large Language Models

### v1

@@ 1 Introduction
With the advent of Large Language Models (LLMs) exhibiting surprising abilities across a variety of language tasks Sakaguchi et al. (2019); Cobbe et al. (2021); Lin et al. (2022); Hendrycks et al. (2020); Zellers et al. (2019); Clark et al. (2018); Gao et al. (2021); Beeching et al. (2023), open-source language models have surged in popularity as part of broader efforts to democratize their accessibility Taori et al. (2023). Furthermore, performance can be improved with custom data that is smaller than their pretraining data, which is known as supervised fine-tuning. Liu et al. (2024b) One type of fine-tuning important to the open-source community is instruction tuning, which allows models to follow a broad or specific-set of instructions and to discuss with users in natural dialogue. Zhang et al. (2023) However, instruction-tuning of open-source LLMs remains limited in multiple areas.
One prominent limitation is the expensive cost of fine-tuning such models, caused by the large number of parameters and data required  Liu et al. (2024a). The popular approach is to scale-up the parameter count and the dataset size, increasing the compute required to train the models. This method to improve LLMs does not lend-well to organizations without massive computing resources, such as the majority of entities who rely on open-source. Another limitation is evaluation, especially for instruction-following models. Evaluation is challenging as a result of LLMs’ inherent open-ended generation, where a space of ideal outputs exists, and existing evaluations usually measure ability in a specific area, such as summarization. Furthermore, evaluation sets may be hard to construct for domain-specific abilities. These two issues limit the continued adoption of open-source models.
@@
By focusing on the problem of coreset selection, we instead take a data-centric approach using influence functions. Influence functions are a method from robust statistics Hampel (1974) and first adapted to neural-network-based machine learning models by Koh and Liang (2017). They approximate how much the model’s prediction changes when a particular training point is removed from the training process via the model gradients.
In this work, we aim to show that Influence Functions are versatile statistical tools that can address two questions on data suitability: 1) What training points are suitable for the test set? and 2) What test points are suitable given that the model was already trained on a particular training set? We use influence functions to identify influential data to address the two issues aforementioned through making fine-tuning more efficient by using less data, and identifying test points that are well-covered by the fine-tuned model.
@@ 2.1 Influence Functions
Influnce functions, a subset of Data Attribution methods, seek to measure the effect of a given training point/s on a trained model. For a given training point, they seek to capture the change in behavior of a trained statistical model had that single training point not been part of the training dataset (leave-one-out-retraining). It outputs a value, called the influence value, for each training point in question. For further discussion on Data Attribution and Influence Functions in machine learning, we refer to Hammoudeh and Lowd (2024).
@@ 2.2 Coreset Selection for LLMs
Coreset selection aims to select a subset of the full training data, such that a model well-fitted to the coreset also fits to the full data. One of the first to study it for big data viewed it as a compression method for large datasets. Phillips (2017) In machine learning, Mirzasoleiman et al. (2020) introduced a coreset algorithm to boost iterative gradient-based training. Coreset selection has been recently explored for LLMs, given the obvious problem of expensive training Li et al. (2023c); Chen et al. (2023) or to understand model properties such as in-context learning Han et al. (2023). Another line of work such as Han et al. (2023) and Wang et al. (2023) apply coreset selection for selecting the pretraining data of language models. Our work differs by focusing on the fine-tuning data, which is relevant to practitioners and to the open-source community.
@@ 3.3 Experimental Setup
For simplicity, we refer to mean perplexity of the model on the validation dataset as perplexity and BERT Score between each validation point and the training dataset as similarity.
@@ 4.1 Results
In Figure 5, we compare two different notions of measuring the importance of the entire training set to each evaluation point: Semantic similarity via BERT scores Zhang et al. (2019) and Influence via influence functions from DataInf. The BERT score of two points is their cosine similarity in the contextual embedding space of BERT Devlin et al. (2018). BERT score aims to capture the semantic similarity between two different sequences of texts. In our experiments, we use DistilBERT Sanh et al. (2019) for efficiency. For this case, we calculate the BERT score of each training point to each test point. Then for each test point, we take the mean of the scores of all training points on that test point. We define this as the similarity score of the entire training set for that particular test point. Therefore for each test point, we obtain a scalar value as the similarity score. Meanwhile, to calculate the influence of the entire training set on each test point, we take the mean of the training point gradients to reduce it into a single point. We then calculate influence values with respect to individual test points as per normal.
For both types of importance, we sort each test point into ascending order and plot the rank of each point. We compare these ranks to the perplexity of the model on the test point, where the model is the Gemma-2B fine-tuned on the entire training set. The red line for each graph denotes the linear regression line. Coefficient values are within a 95% confidence interval, but the intervals are too narrow to visualize. For semantic similarity as a metric, we hypothesize perplexity and semantic similarity to have a negative linear correlation (i.e. a downward trend) because test points with lower semantic similarity to the training set signifies that the test points are from a different distribution than the training set. However, while we observe this downward trend, it is a very weak correlation (coefficient = -0.087), suggesting that semantic similarity as measured by BERT scores is not a strong signal to indicate the fit of the training set to the test points.
In contrast, perplexity and influence values exhibit a stronger correlation. For Influence as a metric, we hypothesise perplexity and Influence to have a positive linear correlation (i.e. an upward trend) because test points assigned negative influence values imply that the training set is a proponent of those test points, and the training set is an opponent for the reverse case. We see this relationship between perplexity and influence, and observe that the correlation is stronger (coefficient = 0.56). This suggests that using influence values provide a stronger signal to indicate whether the model can generalize to a particular test point. Furthermore, using influence values is cheaper compared to using BERT scores, because we can reduce the training set to a single point when representing it as a gradient, compared to calculating point-wise similarities with BERT Scores.
@@ 5 Conclusion
LLM Evaluation There is ongoing work to evaluate instruction-following models on a variety of tasks Beeching et al. (2023); Gao et al. (2021); Hendrycks et al. (2020); Liang et al. (2022), but reaching a consensus on how to best evaluate such general-purpose models is still an open issue. Furthermore, there is an existing line of work that uses a larger model for automatic evaluation Li et al. (2023a); Zhao et al. (2023); Chen et al. (2023); Chiang et al. (2023), most popularly GPT-4 Achiam et al. (2023). However, the decision of the larger model can be biased to models from the same family. Li et al. (2023a) Furthermore, the output is opaque and uninterpretable because there is no guarantee that the accompanying explanation to its answer is faithful. Jacovi and Goldberg (2020)
Group influence
Our formulation of influence looks at the contribution of each points individually, but not their contribution as a group of points. We assumed that removing / adding groups of points based on their individual influences would create the expected effect (e.g. a sufficient learned model). Although we demonstrated this empirically in Section 3.4, we leave to future work the theoretical analysis if there are better methods to select points based on their individual or group influences.
As for selecting a group of points based on their group influences, we would like to stress that our algorithm could be applied for group influence by replacing DataInf with the group influence algorithm of choice. Calculating group influence with influence functions has first been studied by Koh et al. (2019) and algorithms to calculate it are becoming more efficient. Basu et al. (2020)
Average gradient of the tokens as the gradient of the entire sequence
We take the average gradient of the tokens when calculating the gradient of each point to make influence value calculations tractable. Xia et al. (2024) showed that this method penalizes the influence of points with long sequences of text. This is a fundamental limitation for definitions of influence that involve calculating the gradient of sequences of varying lengths. See Table 7 for training examples selected by In2Core.
@@ 6 Ethics Statement
Training LLMs is compute-intensive and thus carbon emissions-heavy. Rillig et al. (2023) Our paper introduces a method to make LLM fine-tuning more efficient, which can lead to positive environmental impact. Using less data to fine-tune translates to less electricity and water consumption, all else being equal. We intend our method to be adopted by practitioners to reduce the cost of training. Our results are also transparent and derived from open-source LLMs & datasets.

### v2

@@ 1 Introduction
With the advent of Large Language Models (LLMs) exhibiting surprising abilities across a variety of language tasks Clark et al. (2018); Zellers et al. (2019); Sakaguchi et al. (2019); Hendrycks et al. (2020); Cobbe et al. (2021); Gao et al. (2021); Lin et al. (2022); Beeching et al. (2023), open-source language models have surged in popularity as part of broader efforts to democratize their accessibility Taori et al. (2023). Furthermore, performance can be improved with custom data that is smaller than their pretraining data, which is known as supervised fine-tuning Liu et al. (2024b). One type of fine-tuning important to the open-source community is instruction tuning, which allows models to follow a broad or specific-set of instructions and to discuss with users in natural dialogue Zhang et al. (2023). However, instruction-tuning of open-source LLMs remains limited in multiple areas.
One prominent limitation is the expensive cost of fine-tuning such models, caused by a large number of parameters and data required  Liu et al. (2024a). The popular approach is to scale-up the parameter count and the dataset size, increasing the computation required to train the models. This method to improve LLMs does not lend-well to organizations without massive computing resources, such as the majority of entities who rely on open-source. Another limitation is evaluation, especially for instruction-following models. Evaluation is challenging as a result of LLMs’ inherent open-ended generation, where a space of ideal outputs exists, and existing evaluations usually measure ability in a specific area, such as summarization. Furthermore, evaluation sets may be hard to construct for domain-specific abilities. These two issues limit the continued adoption of open-source models.
@@
By focusing on the problem of coreset selection, we instead take a data-centric approach using influence functions. Influence functions are a method from robust statistics Hampel (1974) and are first adapted to neural-network-based machine learning models by Koh and Liang (2017). They approximate how much the model’s prediction changes when a particular training point is removed from the training process via the model gradients.
In this work, we aim to show that Influence Functions are versatile statistical tools that can address two questions on data suitability: 1) What training points are suitable for the test set? and 2) What test points are suitable given that the model was already trained on a particular training set? We use influence functions to identify influential data to address the two aforementioned issues by making fine-tuning more efficient by using less data and identifying test points well-covered by the fine-tuned model.
@@ 2.1 Influence Functions
Influence functions, a subset of Data Attribution methods, seek to measure the effect of a given training point/s on a trained model. For a given training point, they seek to capture the change in behavior of a trained statistical model had that single training point not been part of the training dataset (leave-one-out-retraining). It outputs a value, called the influence value, for each training point in question. For further discussion on Data Attribution and Influence Functions in machine learning, we refer to Hammoudeh and Lowd (2024).
@@ 2.2 Coreset Selection for LLMs
Coreset selection aims to select a subset of the full training data, such that a model well-fitted to the coreset also fits to the full data. One of the first to study it for big data viewed it as a compression method for large datasets Phillips (2017). In machine learning, Mirzasoleiman et al. (2020) introduced a coreset algorithm to boost iterative gradient-based training. Coreset selection has been recently explored for LLMs, given the obvious problem of expensive training Li et al. (2023c); Chen et al. (2023) or to understand model properties such as in-context learning Han et al. (2023). Another line of work such as Han et al. (2023) and Wang et al. (2023) apply coreset selection for selecting the pretraining data of language models. Our work differs by focusing on the fine-tuning data, which is relevant to practitioners and to the open-source community.
@@ 3.3 Experimental Setup
For simplicity, we refer to the mean perplexity of the model on the validation dataset as perplexity and BERTScore between each validation point and the training dataset as similarity.
@@ 4.1 Results
In Figure 5, we compare two different notions of measuring the importance of the entire training set to each evaluation point: Semantic similarity via BERTScore Zhang et al. (2019) and Influence via influence functions from DataInf. The BERTScore of two points is their cosine similarity in the contextual embedding space of BERT Devlin et al. (2018). BERTScore aims to capture the semantic similarity between two different sequences of texts. In our experiments, we use DistilBERT Sanh et al. (2019) for efficiency. For this case, we calculate the BERTScore of each training point to each test point. Then for each test point, we take the mean of the scores of all training points on that test point. We define this as the similarity score of the entire training set for that particular test point. Therefore for each test point, we obtain a scalar value as the similarity score. Meanwhile, to calculate the influence of the entire training set on each test point, we take the mean of the training point gradients to reduce it into a single point. We then calculate influence values with respect to individual test points as per normal.
For both types of importance, we sort each test point into ascending order and plot the rank of each point. We compare these ranks to the perplexity of the model on the test point, where the model is the Gemma-2B fine-tuned on the entire training set. The red line for each graph denotes the linear regression line. Coefficient values are within a 95% confidence interval, but the intervals are too narrow to visualize. For semantic similarity as a metric, we hypothesize perplexity and semantic similarity to have a negative linear correlation (i.e. a downward trend) because test points with lower semantic similarity to the training set signifies that the test points are from a different distribution than the training set. However, while we observe this downward trend, it is a very weak correlation (coefficient = -0.087), suggesting that semantic similarity as measured by BERTScore is not a strong signal to indicate the fit of the training set to the test points.
In contrast, perplexity and influence values exhibit a stronger correlation. For Influence as a metric, we hypothesise perplexity and Influence to have a positive linear correlation (i.e. an upward trend) because test points assigned negative influence values imply that the training set is a proponent of those test points, and the training set is an opponent for the reverse case. We see this relationship between perplexity and influence, and observe that the correlation is stronger (coefficient = 0.56). This suggests that using influence values provide a stronger signal to indicate whether the model can generalize to a particular test point. Furthermore, using influence values is cheaper compared to using BERTScore, because we can reduce the training set to a single point when representing it as a gradient, compared to calculating point-wise similarities with BERTScore.
@@ 5 Conclusion
LLM Evaluation There is ongoing work to evaluate instruction-following models on a variety of tasks Hendrycks et al. (2020); Gao et al. (2021); Liang et al. (2022); Beeching et al. (2023); Wang et al. (2024a, b), but reaching a consensus on how to best evaluate such general-purpose models is still an open issue. Furthermore, there is an existing line of work that uses a larger model for automatic evaluation Li et al. (2023a); Chiang et al. (2023); Zhao et al. (2023); Chen et al. (2023); Wang et al. (2024c), most popularly GPT-4 Achiam et al. (2023). However, the decision of the larger model can be biased to models from the same family. Li et al. (2023a) Furthermore, the output is opaque and uninterpretable because there is no guarantee that the accompanying explanation to its answer is faithful Jacovi and Goldberg (2020).
Group influence. Our formulation of influence looks at the contribution of each points individually, but not their contribution as a group of points. We assumed that removing / adding groups of points based on their individual influences would create the expected effect (e.g. a sufficient learned model). Although we demonstrated this empirically in Section 3.4, we leave to future work the theoretical analysis if there are better methods to select points based on their individual or group influences.
As for selecting a group of points based on their group influences, we would like to stress that our algorithm could be applied for group influence by replacing DataInf with the group influence algorithm of choice. Calculating group influence with influence functions has first been studied by Koh et al. (2019) and algorithms to calculate it are becoming more efficient Basu et al. (2020).
Average gradient of the tokens as the gradient of the entire sequence. We take the average gradient of the tokens when calculating the gradient of each point to make influence value calculations tractable. Xia et al. (2024) showed that this method penalizes the influence of points with long sequences of text. This is a fundamental limitation for definitions of influence that involve calculating the gradient of sequences of varying lengths. See Table 7 for training examples selected by In2Core.
@@ 6 Ethics Statement
Training LLMs is compute-intensive and thus carbon emissions-heavy Rillig et al. (2023). Our paper introduces a method to make LLM fine-tuning more efficient, which can lead to positive environmental impact. Using less data to fine-tune translates to less electricity and water consumption, all else being equal. We intend our method to be adopted by practitioners to reduce the cost of training. Our results are also transparent and derived from open-source LLMs & datasets.